OpenStudy (anonymous):

In how many different ways can you answer a multiple choice test that has 5 questions and 3 choices for each answer?

OpenStudy (anonymous):

125 243

OpenStudy (kropot72):

Each of the 3 possible answers to the first question can be combined with any of the 3 possible answers to the second question. Therefore the number of ways of answering the first two questions = 3 * 3 = 9. Each of the 3 possible answers to the third question can be combined with any of the 9 possible ways of answering the first two questions. Therefore the number of ways of answering the first three questions = 3 * 3 * 3 = 27. Carrying on the reasoning as above, the number of ways of answering the five questions = 3 * 3 * 3 * 3 * 3 =?